Braxton Enterprises currently has debt outstanding of $ 35 million and an interest rate of 9 % Braxton plans to reduce its debt by repaying $ 7 million in principal at the end of each year for the next five years. If​ Braxton's marginal corporate tax rate is 40 %, what is the interest tax shield from​ Braxton's debt in each of the next five​ years?

The interest tax shield in year one is -

the interest tax shield in year two (round to three decimal place)

the interest tax shield in year three (round to three decimal place)

the interest tax shield in year four( round to three decimal place)

the interest tax shield in year five (round to three decimal place)
Answer #1
i ii=i*9% iii iv=i-iii v=ii*40%
year Beginning loan balance Interest amount Principal repayment ending loan balance Interest tax shield
1 35000000 3150000 7000000 28000000 1260000
2 28000000 2520000 7000000 21000000 1008000
3 21000000 1890000 7000000 14000000 756000
4 14000000 1260000 7000000 7000000 504000
5 7000000 630000 7000000 0 252000